# EpidemIQs: Prompt-to-Paper LLM Agents for Epidemic Modeling and Analysis

### v1

@@ III-C Memory
Two specialized memory structures are considered for the agents. The scientist agent’s memory consists of short-term and long-term memory. The short-term memory encompasses all the current conversation details, including role description, input query, agent’s internally generated responses, interaction with tools, conversations with other agents, and the final output. The long-term memory stores all previous conversation history in a database, enabling semantic retrieval based on the current query.
III-D Structured Outputs
@@
Typically, LLMs generate free-format texts as a natural language output, which does not follow a strict structure. However, the performance of the collaborative LLM agents requires precise and careful interoperability, as information sharing and communication can significantly impact overall outcomes [44]. Therefore, to make the data transition between each phase of the process robust, reliable, and more predictable, there are constraints defined for each team output structure, forcing them to produce output in a predefined structure (e.g., JSON schema). This not only facilitates data retrieval and improves prompt development efficiency, but it also enables automatic validation of each team’s output [56], ensuring all required fields are generated accordingly, and there is no missing data. If not, the agent retries until it fulfills the task or reaches the maximum number of tries allowed. While the output structure of each scientist contains different fields, all scientists share one common element: the reasoning and logical justification behind their choices and results. Each scientist is expected to defend their decisions against hypothesized criticisms, demonstrating why these were the most appropriate outcomes.
III-E Autonomous and Copilot Operation Modes
@@
The EpidemIQs can perform in two modes: end-to-end autonomous and cooperative modes. The latter is named as a copilot, which can operate as an assistant to the user. The autonomous mode does not require human intervention, but rather initially provides the query. The copilot mode enables the human to intervene and provides feedback and instruction on the agent’s output, or asks for more actions to guide the agent toward the desired direction. After each phase, the human is asked to review the results, and they can either mark them as complete and proceed to the next phase or ask for new actions by providing comments.
III-F Implementation
@@
The framework is implemented in Python using the Pydantic AI framework. Autonomous data acquisition is supported through API integrations using the Semantic Scholar API for literature retrieval and the Tavily API for web-scale information access. For multimodal processing, a Vision-Language Model (VLM) is implemented via the OpenAI SDK with Base64 decoding to enable visual data interpretation. Short-term memory utilizes the LLM’s context window, while long-term memory is stored in JSON and managed through Pydantic AI memory functions. Finally, we used Pydantic data models for schema validation and structural consistency across agent interactions.
@@ IV Evaluation Approach
To assess the performance of the framework, we designed five questions, shown in Figure 4, with increasing depth and complexity.
@@
The first question is the simplest, requiring an analytical investigation and comparison of stochastic simulations on heterogeneous versus homogeneous networks to assess how structural differences shape disease spread. The second focuses on identifying why the chain of transmission breaks and validating hypotheses through stochastic modeling. The third resembles the first but introduces a temporal network, which exceeds the scope of the provided simulation tools, testing system adaptability.
@@
The fourth demands both analytical reasoning (or knowledge from literature) and an extension of the SIS model to a competitive $SI_{1}I_{2}S$ form on a multiplex network, exploring conditions for coexistence. The fifth is the most challenging, requiring analytical evaluation of random versus targeted vaccination strategies. Here, the key challenge is to notice the implicit constraint to ensure that the contact network contains enough degree-10 nodes to make targeted vaccination feasible while preserving its structural properties.
@@
Collectively, these questions span the key aspects of network epidemics, such as topology, temporal structure, multilayer interactions and competing pathogens, termination mechanisms, network-aware interventions, physics-aware contact network, etc, providing a holistic evaluation framework. Only the first two questions were involved in the test and trials through the design of the EpidemIQs orchestration, and the last three were unknown to the framework.
Epidemiological research is inherently stochastic and complex, and admits multiple valid modeling approaches, leading to the absence of a unified standardized benchmark for evaluating the end-to-end epidemic research pipeline. Therefore, to evaluate our framework, we adopted an assessment strategy aligned with how the scientific community assesses research manuscripts through the peer review process. Consequently, we considered:
  • Human evaluation by peer review of written outputs and appraisal of deliverables. Specifically, we asked five blinded domain experts to review the manuscripts generated by EpidemIQs, based on the criteria in Supplementary Table S7.
@@
  • LLM-as-Judge evaluation using GPT-4o to evaluate generated papers autonomously based on the same assessment criteria as human reviewers to deliver timely evaluation.
@@
  • Completion success rate defined as the proportion of trials in which EpidemIQs autonomously executed the full workflow and produced a scientific report containing all pre-specified sections, independent of content quality.
  • Computational costs in terms of turnaround time, number of generated tokens, and total API cost.
Together, these measures capture both the scientific quality of the outputs and the reliability of the autonomous workflow. We compared our framework with a multi-turn single agent, such that an agent plans its next steps autonomously for up to the maximum number of steps, $s_{max}$. The single agent is instructed to perform all five phases with clear instructions (same as the agents in the EpidemIQs) and the same tools as in the multi-agent framework, along with VisionExpert to enable the agent to analyze the figures. This single agent needs to have an LLM with a long context window; therefore, we used OpenAI gpt-4.1-2025-04-14, as it has reasoning capability and support of structured output with a context window of about one million tokens, along with o3-2025-04014 as the current highly capable OpenAI model with a 200K tokens context window.
To prevent the prompt drift, in each turn we re-inject the main task into memory as **Reminder** main query is {query}. Do NOT lose sight of this, ensuring that the agent’s working context always contains the overarching objective. We considered the single agent as a baseline, and compared it with EpidemIQs’ implementation based on computational statistics, performance analysis, and review scores of the generated papers.
@@ V Results and Discussion
Across the five epidemic questions, EpidemIQs completed every phase of the pipeline as Discovery, Modeling, Simulation, Analysis, and Report writing in autonomous mode, achieving a 100% completion success rate in generating the final report file. Automated rubric scoring by the LLM‑as‑Judge assigned a mean quality score of 9.04±0.21 out of 10, while four network science Ph.D. students and one faculty reviewer gave an average expert rating of 7.98±0.35, praising the reports’ methodological soundness, clarity, and depth. Below, we summarize how the framework navigated each question, highlighting key decisions and results. The generated papers and a more comprehensive analysis are provided in the supplementary material.
Figure 5: Examples of paper generated by EpidemIQs for the second question.
Figure 6: An example of reasoning information NetworkScientist generated for question one
Question 1. In the first phase, analytical percolation theory and the literature confirm that outbreak behavior can differ according to the degree of heterogeneity. They suggest that, for a heterogeneous network, a scale-free graph, whose node-degree distribution follows a power law, producing a few highly connected hubs among many sparsely linked nodes, is an appropriate candidate to compare with a well-mixed homogeneous network, in which every node is connected to every other node, resulting in a fully connected graph. Because the question does not specify the intrinsic reproduction number, $R_{0}$, a value of 2.5 is adopted to reflect a disease that can be characterized by an SEIR model, such as COVID-19, for which $2<R_{0}<3$. Consequently, the deterministic model parameters are chosen to reflect typical values observed in COVID-19, incubation rate as $\sigma_{E\rightarrow I}=0.2/day$ (average incubation period of 5 days), recovery rate, $\gamma_{I\rightarrow R}=0.1/day$ (average infectious period of 10 days), and transmission rate $\beta_{S\rightarrow E}=0.25$. The framework then generates both graphs, initializes them with the five infected nodes, and tunes the transmission rate of the SEIR model on the basis of the intrinsic $R_{0}$ and a mean-field approximation to consider the effect of different network topologies.
In the simulation phase, three scenarios are executed for 300 stochastic realizations each: (i) random infectious seeding in the homogeneous network, (ii) random infectious seeding in the scale-free network, and (iii) infectious seeding at the highest-degree hubs in the scale-free network. Analysis of the simulation results corroborates the analytical theory. The fully connected homogeneous network exhibits a rapid, high-magnitude outbreak with near-complete infection penetration. By contrast, the scale-free network shows markedly different epidemic behaviour in both seeding schemes, with a much later epidemic peak, a lower peak prevalence, and a substantially smaller final epidemic size ($\approx 30\%$). Notably, hub-based infection seeding accelerates epidemic progression and yields a higher average final epidemic size, confirming the role of superspreaders in triggering more rapid outbreaks in networks with heavy-tailed degree distributions. Hence, the stochastic results confirm that degree heterogeneity decisively reshapes outbreak timing and magnitude. Complete description and results for this question are provided at Supplementary Section S1.1.
Figure 6 illustrates NetworkScientist’s reasoning and logic, exemplifying how each scientist justifies its task completion and the rationale behind its choices, thereby enhancing the interpretability of the outcomes.
@@
Figure 7: An overview of the workflow for question four, highlighting each step’s goals, main actions, and outputs. The results have been truncated for clearer visualization. At the bottom, plots generated during the modeling and simulation phases are provided.
Question 2. Starting from the discovery phase, analytical results showed that the chain of transmission breaks when the effective reproduction number drops below unity, $R_{e}<1$. This can occur either when the susceptible population falls beneath a critical threshold or when the transmissibility of infectives decreases. Moreover, contact patterns (e.g., network topology) influence whether an outbreak sputters or ignites, thereby shaping its overall dynamics. To validate these findings, consistent with the surveyed literature, the framework employed an SIR model with two basic reproduction numbers: a supercritical case, $R_{0}=3$, and a subcritical case, $R_{0}=0.5$.
To isolate structural effects, two networks with identical average degree $\langle k\rangle=10$ and N=1, [^0] N=1{,}000 nodes were generated. The first is an Erdős–Rényi (ER) network representing homogeneous mixing, where each pair of nodes connects with probability $p$. The second is a scale-free Barabási–Albert (BA) network capturing heterogeneity; after an initial seed, each incoming node forms $m$ links to existing nodes, producing a heavy-tailed degree distribution.
Simulations began with ten randomly seeded infectives. For each network and $R_{0}$ value, transmission rates were adjusted (via mean-field approximation) to account for degree heterogeneity, and 75 stochastic realizations of each scenario were executed. In the supercritical scenario ($R_{0}=3$), the epidemic peaked rapidly and faded as susceptibles were depleted. In the ER network, the average susceptible count fell from 990 to $\approx 115$ and the recovered count rose to $\approx 885$, indicating herd-immunity termination. The BA network also exhibited fade-out but progressed faster and left a larger susceptible pool (final average epidemic size $\approx 520$, remaining susceptibles $\approx 480$), reflecting heterogeneity-mediated herd immunity.
For the subcritical scenario ($R_{0}=0.5$), intrinsic transmission inefficiency halted spread in both network topologies. Overall, the simulations confirm the analytical prediction that complete exhaustion of susceptibles is unnecessary; herd immunity can emerge when the susceptible fraction drops below a critical level or when per-contact transmissibility is sufficiently low. Figure 5 shows an overview of the generated paper, and details of the framework’s workflow are provided in the Supplementary Section S1.2.
Question 3. In the discovery phase, statistical analysis using a mean-field approximation derived expressions for the epidemic threshold in an activity-driven network, where each node activates at rate $\alpha=0.1$ and, upon activation, instantaneously forms $m$ edges, which are then removed in the next step. Because neither $m$ nor the recovery rate was specified, the experiment adopted a unit recovery rate and set $m=5$ as a plausible value. The analytical results indicated that, for $R_{0}=3$, a final epidemic size of 94% should be expected under these parameters.
To compare the analytical solution for the temporal network having 1000 nodes with its time-aggregated static counterpart, a static network was created by aggregating all temporal contacts over $T=1000$ time steps; for every node pair $(i,j)$, the edge weight equaled the total number of contacts during $T$. The recovery rate of one was chosen, and the infection rate was then calibrated for each network-driven SIR simulation so that they matched the mean-field derivations and were consistent with the respective network context (temporal or static). One random infectious node was selected as the initial condition, with all other nodes susceptible.
One hundred stochastic realizations were performed for each scenario, using FastGEMF for the time-aggregated static network and a custom discrete-time Markov-chain engine for the temporal network. The results revealed sharp contrasts in epidemic size and speed between the temporal and static scenarios. The temporal activity-driven network produced significantly smaller and slower epidemics relative to both the static-network simulation and the analytic predictions based on aggregated parameters. Specifically, the temporal SIR process yielded a mean final epidemic size of approximately 20.5% with a large standard deviation of 28.9%, indicating high stochastic variability and frequent epidemic die-out events at the chosen parameters. Such early stochastic die-out is ignored in mean-field ODEs, which assume deterministic branching of infections and never allow the process to fizzle when the first few seeds recover before transmitting the pathogen.
By contrast, the static-network simulation produced near-complete infection of the population (mean final size $\approx 81\%$) in a rapid fashion, closely approaching the analytical prediction of 94.05% derived from the classic mean-field equation. These results highlight the profound mechanistic effect of temporal causality on disease-transmission pathways. In the temporal network, edges represent instantaneous contacts that exist only transiently and are reorganized stochastically at each time step. Consequently, the temporal ordering of contacts restricts the accessibility of transmission chains, effectively increasing the epidemic threshold and reducing the reachable susceptible population. Many potential transmission paths present in the aggregated static network are simply infeasible in the temporal network because the required contact timings do not overlap to support causally coherent infections. Supplementary Section S1.3. provides the detailed outcomes for this question.
Figure 8: An overview of the workflow for question five, outlining each step’s goals, main actions, and outputs. For clarity, the results have been truncated. At the bottom, plots display the degree distribution, highlighting the large number of nodes with degree 10 and confirming that enough nodes are available for targeted vaccination, together with simulation results for four cases: no vaccination (baseline), vaccination below the critical threshold, at the threshold, and above the threshold.
Figures 7 and 8 display the results and process pipelines for questions 4 and 5, respectively. For the question 4 scenario, the framework could successfully find the criteria for the coexistence of the competitive exclusive viruses over multilayer networks, extend the analytical mean-field point of view that exists in literature, and show the dominance, coexistence, and extinction regimes in stochastic processes by designing the network properties that allow coexistence and tuning the parameters. Finally, for the last question, it not only determined the vaccination threshold for random and targeted cases but also carefully modeled the contact network to maintain the aforementioned centralities while having enough nodes to verify the targeted vaccination threshold. The presented results are truncated, and please refer to Supplementary Sections S1.4 and S1.5. for full detailed results of questions 4 and 5, respectively.

### v2

@@ III-C Memory
Two specialized memory structures are considered for the agents. The scientist agent’s memory consists of short-term and long-term memory. The short-term memory encompasses all the current conversation details, including role description, input query, agent’s internally generated responses, interaction with tools, conversations with other agents, and the final output. The long-term memory stores all previous conversation history in a database, enabling memory retrieval based on the current query.
III-D Structured Outputs and Operation Modes
@@
Typically, LLMs generate free-format texts as a natural language output, which does not follow a strict structure. However, the performance of the collaborative LLM agents requires precise and careful interoperability, as information sharing and communication can significantly impact overall outcomes [37]. Therefore, to make the data transition between each phase of the process robust, reliable, and more predictable, there are constraints defined for each team output structure, forcing them to produce output in a predefined structure (e.g., JSON schema). This not only facilitates data retrieval and improves prompt development efficiency, but it also enables automatic validation of each team’s output [45], ensuring all required fields are generated accordingly. If not, the agent retries until it fulfills the task or reaches the maximum number of tries allowed. While the output structure of each scientist agent contains different fields, they all share a common element: the reasoning and logical justification behind their choices and results. Each scientist agent is expected to defend their decisions against hypothesized criticisms, demonstrating why these were the most appropriate outcomes.
@@
Furthermore, EpidemIQs can operate in two modes: end-to-end autonomous and cooperative. The latter is named as a copilot, which needs the user to be actively in the loop to proceed. The autonomous mode does not require human intervention during pipeline execution; instead, it initially provides the query. The copilot mode enables the human to intervene, providing feedback and instruction on the agent’s output or asking for additional actions to guide the agent toward the desired direction. After each phase, the human is asked to review the results, and they can either mark them as complete and proceed to the next phase or ask for new actions by providing comments.
III-E Implementation
@@
The framework is implemented in Python by the Pydantic AI framework. Autonomous data acquisition is supported through API integrations using the Semantic Scholar API for literature retrieval and the Tavily API for web-scale information access. Vision-Language Model (VLM) is implemented via the OpenAI SDK with BinaryContent format to enable efficient visual data interpretation for multimodal processing. Short-term memory utilizes the LLM’s context window, while long-term memory is stored in JSON and managed through Pydantic AI memory functions. Finally, we used Pydantic data models for structured outputs for schema validation. For complete implementation details, including pinned software versions, random seed configurations, and a manifest of all experimental artifacts, refer to Supplementary Section LABEL:S-sec:repro-summary.
@@ IV Evaluation Approach
To demonstrate how the framework addresses various tasks, we adopted five epidemic questions (Figure 4), with increasing depth and complexity, to cover different aspects of epidemic modeling.
@@
The first question in Figure 4 is the simplest. It requires the framework to analytically investigate the individual behavior effect on spread, design simulation setup by creating the SEIR model and at least two contact network structures, one heterogeneous and one homogeneous, to compare the stochastic results. The goal is to analyze how structural heterogeneity in host contact patterns shapes population level disease dynamics.
@@
The second focuses on identifying why the chain of transmission breaks and validating hypotheses through stochastic modeling [43]. The third scenario introduces a temporal contact pattern, which exceeds the scope of the provided simulation tools and instruction, testing system adaptability[46]
@@
The fourth demands both analytical reasoning (or knowledge from literature [47]), an extension of the SIS model to a competitive $SI_{1}I_{2}S$ form on a multiplex network, exploring conditions for coexistence, and simulation on multilayer networks, for which no instructions were provided regarding the multilayer networks setup. The fifth is the most challenging, requiring analytical evaluation of random versus targeted vaccination strategies. Here, the key challenge is to notice the implicit constraint to ensure that the contact network contains enough degree-10 nodes to make targeted vaccination feasible while preserving its structural properties (January 2025 question of the month from Network Science Society).
@@
Collectively, these questions span the key aspects of network epidemics, such as topology, temporal structure, multilayer interactions and competing pathogens, termination mechanisms, network-aware interventions, physics-aware contact network, etc, providing a holistic evaluation framework. We emphasize that these scenarios function as architectural stress tests for the autonomous research pipeline; they are intended to demonstrate the framework’s methodological coherence, distinct from the empirical calibration required for real-world epidemiological forecasting.Only the first two questions were involved in the test and trials through the design of the EpidemIQs orchestration, and the last three were unknown to the framework and were out of the immediate scope of the tools or system prompts provided to the system.
Having a standardized benchmark for evaluating an agentic AI system for science remains an open and non-trivial challenge. Unlike coding, retrieval, or mathematical tasks with binary success conditions, scientific research is inherently open-ended. Recent works have adopted varying assessment strategies. For example, focusing on the overall usefulness of the results [27, 48], relying exclusively on AI-based judgment [31], labeling randomly sampled statements from reports as correct or incorrect[49], or highlighting successful outcomes in specific case studies to demonstrate applicability [23, 28, 29]. In epidemiology, this challenge is compounded by the stochastic nature of the problem and the existence of multiple valid modeling approaches for a single query. Therefore, there is no single ground truth against which to measure performance mechanically. Consequently, we adopted a multi-faceted assessment strategy aligned with the scientific peer-review process, evaluating both technical validity and report quality. We considered the following metrics:
  • Human Expert Evaluation to measure overall correctness and quality. We employed blinded peer review by domain experts to assess the generated manuscripts based on the criteria in Supplementary Table LABEL:S-tab:evaluation_criteria. They also inspected the generated artifacts (code, plots, and simulation data) to verify Technical Soundness and Experimental Rigor, ensuring the reported results were supported by the underlying work and that they could obtain consistent results through the same random seeding of executed experiments. They also evaluated the quality of the reports and their relevance and alignment with results.
@@
  • LLM-as-Judge Evaluation using GPT-4o to autonomously evaluate the generated papers using the same assessment criteria as human reviewers, providing a scalable metric for relative quality comparison.
@@
  • Workflow Completion Rate is defined as the proportion of trials in which EpidemIQs autonomously executed the full workflow and produced a scientific report containing all pre-specified sections, independent of content quality.
  • Task Success Rate, where success is defined as a binary metric evaluated by human experts based on two criteria: (i) the scientific correctness of the conclusion, and (ii) whether the conclusion is substantiated by the stochastic simulation results. This metric provides a feasible approach to analyze the approximate usefulness and correctness across all trials.
  • Computational Costs by tracking the turnaround time, total token usage, and API costs to assess the efficiency of the framework.
While these measures provide comparative performance indicators (descriptive quantitative statistics and rubric-based judgments) rather than statistically validated benchmarks, they capture both the scientific quality of the outputs and the reliability of the autonomous workflow. For baseline configuration to evaluate the architectural benefits of the EpidemIQs orchestration, we compared our framework against a robust, multi-turn single-agent baseline. To ensure a fair comparison, this baseline was designed with structural guidance equivalent to EpidemIQs. The single-agent operates iteratively, autonomously planning its next steps up to a maximum limit, $s_{max}$. It is provided with the same curated workflow instructions, phase-specific structured prompts, and definitions of required output fields as the specialized agents in EpidemIQs. However, to strictly avoid penalizing the baseline, we modified the prompts to remove multi-agent specific role designations (e.g., You are a full professor in ...), as these shifting persona constraints within a single context window would risk inducing role confusion or behavioral inconsistency. Instead, the single-agent acts as a unified scientific entity, focusing purely on the procedural instructions and objectives of each phase.
Furthermore, the single-agent is equipped with the identical toolset used in the multi-agent framework. This includes access to Vision Expert, which offloads the token-heavy process of image analysis. To ensure the agent’s capability to handle the load of handling all scientific phases within a single context, we used LLMs with large context capabilities: OpenAI gpt-4.1-2025-04-14 ($\approx$1 million token context window with reasoning capabilities) and o3-2025-04-14 (200K token context window and high reasoning capability). To mitigate context drift as a common issue in long-horizon tasks, we implemented a memory reinforcement mechanism. In each turn, the primary objective is re-injected into the context as **Reminder** main query is {query}. Do NOT lose sight of this. This setup ensures that the baseline represents a strong comparison between monolithic context coherence versus the distributed attention mechanism of the EpidemIQs multi-agent architecture. We provide a comparison between their performance based on computational statistics, performance analysis, and review scores of the generated results.
@@ V Results and Discussion
Across the five epidemic scenarios, EpidemIQs demonstrated high architectural stability, with our evaluation indicating a workflow completion rate of 100% across 100 trials. Automated rubric scoring by the LLM‑as‑Judge assigned a mean quality score of 9.04±0.21 out of 10, while four network science Ph.D. students and one faculty reviewer gave an average expert rating of 7.98±0.35, praising the reports’ methodological soundness, clarity, and depth. The task success rate varied based on the complexity of the task and the alignment of the provided tools. For the first two scenarios (Questions 1 and 2), the framework successfully addressed them in approximately 90% of trials. The temporal contact pattern (Question 3) presented the greater challenge, with a task success rate of 65%. For identifying viral coexistence regions (Question 4), it achieved 75%, and for calculating targeted vaccination thresholds (Question 5), 80% of the time it could calculate and validate vaccination thresholds, resulting in an average of 79±8.60% overall task success rate. In this section, first we briefly discuss the successful results of the framework, followed by an ablation study V-B, limitations V-D, a detailed failure taxonomy V-C1, and finally, ethical concerns are discussed V-E).
Figure 5: An example of reasoning information Network Scientist generated for Question 1. This highlights how the agent justifies the selection of scale-free networks to model heterogeneity against a homogeneous baseline.
Figure 6: An example of paper generated by EpidemIQs for the second question.
Question 1. EpidemIQs successfully validated the fundamental impact of network topology on disease spread dynamics. By modeling two contrasting topologies, a heterogeneous scale-free network and a homogeneous complete graph, the framework demonstrated how degree heterogeneity and superspreader nodes significantly alter epidemic dynamics compared to well-mixed models. The stochastic experiments (300 realizations) and the results analysis confirmed that structural heterogeneity alters the final epidemic size, peak infection size and time, etc., aligning with theoretical expectations. Figure 5 illustrates the Network Scientist’s reasoning, highlighting how the agent justified its modeling choices to capture realistic contact patterns. Full workflow details and the generated report are available in Supplementary Section LABEL:S-sec:q1.
@@
Figure 7: An overview of the workflow for Question 4, highlighting each step’s goals, main actions, and outputs. The results have been truncated for clearer visualization. At the bottom, plots generated during the modeling and simulation phases are provided, showing extinction, coexistence, and dominance regions.
Question 2. The framework correctly determined that transmission chains break when the effective reproduction number drops below unity ($R_{e}<1$), rather than solely due to the complete depletion of susceptibles. To validate this, the system engaged an SIR model across both supercritical ($R_{0}=3$) and subcritical ($R_{0}=0.5$) regimes on Erdős-Rényi and Barabási-Albert networks. The resulting manuscript (visualized in Figure 6) presented a robust analysis showing that outbreaks can self-extinguish due to low transmissibility or network constraints even when a large fraction of the population remains susceptible. This demonstrates the framework’s capacity for validating counterintuitive epidemiological concepts through rigorous simulation, as discussed in detail in Supplementary Section LABEL:S-sec:q2
Question 3. This scenario tested the system’s adaptability by introducing a temporal contact network that is out of the scope of the provided simulation tools. EpidemIQs successfully derived a mean-field approximation for the epidemic threshold. It identified that the temporal ordering of contacts imposes structural constraints that increase this threshold. By comparing the temporal network against a time-aggregated static counterpart (where edge weights represented interaction frequency), the analysis revealed a clear divergence: the temporal network resulted in smaller and slower outbreaks. Critically, the framework demonstrated awareness of its computational constraints, explicitly noting in the final report that the provided engine (FastGEMF) was restricted to static experiments, necessitating a custom simulation algorithm and comparative analytical approach, as detailed in Supplementary Section LABEL:S-sec:q3
Figure 7 displays the results and process pipeline for Question 4. In this scenario, the framework successfully identified the criteria for the coexistence of competitive exclusive viruses over multilayer networks and extended the analytical mean-field approach in the literature. It demonstrated the dominance, coexistence, and extinction regimes in stochastic processes by designing contact network properties that enable coexistence and varying the model parameters. Full detailed results are provided in Supplementary Section LABEL:S-sec:q4.
Finally, for Question 5, the framework not only determined the vaccination thresholds for random and targeted vaccination strategies (75% and 11.25%, respectively), but also carefully modeled a configuration network to preserve the aforementioned centrality measures while ensuring a sufficient number of nodes to verify the targeted vaccination threshold. Stochastic realizations demonstrated the threshold’s effectiveness by examining outcomes in regions both above and below it. Supplementary Figure LABEL:S-fig:workflowQ5 presents the truncated workflow results for Question 5, with additional details provided in Supplementary Section LABEL:S-sec:q5.